# Psi-GAN: A power-spectrum-informed generative adversarial network for the emulation of large-scale structure maps across cosmologies and redshifts

### v2

@@ 3.1 Data generation
Secondly, we observe a discrepancy in the power spectrum of the generated lognormal field and the measured power spectrum from the Quijote slice. This can be attributed to correlations in phases being introduced when converting the Quijote initial conditions (obtained by second-order Lagrangian perturbation theory) to a density field. We correct for this discrepancy by iteratively re-scaling the Gaussian power spectrum by the ratio of the power spectrum of the lognormal field output and the measured power spectrum of the target slice at each $k$. We iterate through this process until the discrepancy is less than $0.1$ per cent at all $k$ values.
@@ 3.2 Model architecture
We train a Wasserstein GAN with gradient penalty (WGAN-GP; Arjovsky et al., 2017; Gulrajani et al., 2017) consisting of a generator and a critic. In a traditional GAN, the generator and critic are adversarially trained in tandem in order to produce generated data that is identical to real data. Our approach builds physics into the critic of the GAN to constrain the generator to produce data that is physically consistent with the target domain. A full schematic of the Psi-GAN framework can be found in Figure 1. In this figure, we demonstrate how an emulation can be generated by feeding a lognormal density field, along with its associated cosmology and redshift, into the generator. This framework is trained via a loss function which depends on the output of our physics-informed critic, which takes as inputs either an emulated or $N$-body map, its associated power spectrum, its associated cosmology and redshift, and finally the power spectrum of the corresponding lognormal density field.
@@
GANs for image synthesis often use a purely feature-based network for the critic, however using a fully-convolutional critic resulted in the generator altering the power spectrum of its input when attempting to generate a more realistic output. The power spectrum path was then added to the critic in order to guide the generator towards not altering the power spectrum. As the lognormal input to the generator has a matching power spectrum to their corresponding $N$-body slices, any deviation away from this would be indicative of a generated map, and therefore this information is extremely helpful for the critic to be able to differentiate between a real and generated image. This information is also able to be backpropagated through the critic and generator networks to ensure that the generator is trained to capture features at all scales.
@@ 3.3 Training
We train using the Adam optimiser (Kingma & Ba, 2014) with a base learning rate $\alpha=10^{-4}$, and decay parameters $\beta_{1},\beta_{2}=0.5,0.9$. Following Heusel et al. (2017), we increase the learning rate for the critic by a factor of $f=3$ while using the base rate for the generator. We allow the model to train for an initial period of three epochs, after which we half the learning rate after every epoch where the validation loss increases. We also employ gradient clipping to clamp the magnitude of the gradients to a maximum value of 1 00010001\,0001 000. The gradient penalty term in the loss function should act to keep gradients close to unity, however there is a warm up period until it is able to have its intended effect. Clipping the gradients was found to be useful in avoiding overflow errors before the gradient penalty took effect.
@@ 3.4.5 Bispectrum
We measure the bispectra and reduced matter bispectra based on an estimator of the binned bispectrum (Coulton et al., 2019; Coulton & Spergel, 2019). We report the measurements as a function of the angle $\theta$ between the vectors $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ using two different configurations: $(k_{1},k_{2})=(0.4,0.4)~{}h~{}\mathrm{Mpc}^{-1}$ and $(k_{1},k_{2})=(0.4,0.6)~{}h~{}\mathrm{Mpc}^{-1}$. Bispectra can be measured along different triangle configurations, and it is important to use many configurations when using the bispectrum as a statistical tool in order to break degeneracies when inferring cosmological parameters (Berge et al., 2010). Therefore we measure the bispectra and reduced bispectra using two configurations: $(k_{1},k_{2})=(0.4,0.6)~{}h~{}\mathrm{Mpc}^{-1}$ and $(k_{1},k_{2})=(0.4,0.4)~{}h~{}\mathrm{Mpc}^{-1}$.
@@ 4.1 Randomised test set
We see agreement to within ${\sim}5$ per cent for all metrics, with the exception of the pixel counts at low values of $\mathrm{ln}(1+\delta)$. This is due to the baseline count for the $N$-body simulation being very low (${\sim}10^{2}$), and thus making the relative differences sensitive to small changes in pixel counts.
@@ 5 Conclusions
The largest shortcoming of Psi-GAN is its slightly weaker performance in constraining the power spectrum when tasked with interpolating between cosmologies. In our tests, the power spectrum of samples generated by Psi-GAN showed showed less agreement with $N$-body simulations when interpolating between the cosmologies used in the Quijote simulations. An approximately ${\sim}50$ per cent greater coverage of cosmology space should be enough to reduce this shortcoming, however this would require significant resources to generate. Another potential method of improving this would be to pre-train the model architecture to be able to reconstruct lognormal random fields across an extensive dataset before fine-tuning the model to transform lognormal random field to more accurate emulations.
We used saliency mapping techniques to investigate further architectural improvements to Psi-GAN, which highlighted a slight weakness in capturing long-range correlations as well as a small issue with pixel-scale artefacting. Increasing the depth of the generator by another step (i.e. including an extra set of down-sample and up-sample blocks) should help Psi-GAN model long-range correlations better as the latent space will be more compressed and information will propagate more efficiently across the simulation box. Adding additional ConvNeXt blocks after the first convolution, and before the last convolution should aid in modelling small scales and reduce artefacting.

### v3

@@ 3.1 Data generation
Secondly, we observe a discrepancy in the power spectrum of the generated lognormal field and the measured power spectrum from the Quijote slice. This can be attributed to correlations in phases being introduced when converting the Quijote initial conditions (obtained by second-order Lagrangian perturbation theory) to a density field. We correct for this discrepancy by iteratively re-scaling $P_{\mathrm{G}}(k)$, which is used to generate the lognormal field $\delta_{\mathrm{LN}}$. Each iteration involves generating a lognormal field from $P_{\mathrm{G}}(k)$ as per Equation 5, measuring its power spectrum $P_{\mathrm{LN}}(k)$, calculating the ratio of $P_{\mathrm{LN}}(k)$ to the target power spectrum $P(k)$ at each $k$, and then rescaling $P_{\mathrm{G}}(k)$ by this ratio at each value of $k$. This process is iterated through until $P_{\mathrm{LN}}(k)$ matches the target power spectrum $P(k)$ to within a $0.1$ per cent discrepancy at all values of $k$.
@@ 3.2 Model architecture
We train a Wasserstein GAN with gradient penalty (WGAN-GP; Arjovsky et al., 2017; Gulrajani et al., 2017) consisting of a generator (${\sim}2.06\times 10^{7}$ parameters) and a critic (${\sim}2.59\times 10^{7}$ parameters). In a traditional GAN, the generator and critic are adversarially trained in tandem in order to produce generated data that is identical to real data. Our approach builds physics into the critic of the GAN to constrain the generator to produce data that is physically consistent with the target domain. A full schematic of the Psi-GAN framework can be found in Figure 1. In this figure, we demonstrate how an emulation can be generated by feeding a lognormal density field, along with its associated cosmology and redshift, into the generator. This framework is trained via a loss function which depends on the output of our physics-informed critic, which takes as inputs either an emulated or $N$-body map, its associated power spectrum, its associated cosmology and redshift, and finally the power spectrum of the corresponding lognormal density field.
@@
GANs for image synthesis often use a purely feature-based network for the critic, however using a fully-convolutional critic resulted in the generator altering the power spectrum of its input when attempting to generate a more realistic output. The power spectrum path was then added to the critic in order to guide the generator towards not altering the power spectrum. As a lognormal input to the generator has a matching power spectrum to its corresponding $N$-body slice, any deviation away from this would be indicative of a generated map. The power spectrum path is constructed such that it calculates an elementwise difference between the lognormal and emulated power spectra. This is then fed into the classification head to aid the critic in differentiating emulated images as any significant deviation from the lognormal power spectrum can be used to easily identify an emulation, thus aiding the critic in achieving its goal. This in turn forces the generator to learn how to maintain the power spectrum of its input so that it can successfully “fool” the critic. This information is also able to be backpropagated through the critic and generator networks to ensure that the generator is trained to capture features at all scales.
This approach was favoured over the more traditional method of adding a power spectrum term directly to the loss function as the loss function was found to be extremely sensitive to the weighting of this additional term, often resulting in non-convergent training. Determining the optimal value of this weighting would require a brute-force hyperparameter search, which would be less efficient than our method of introducing a power spectrum path to the critic and allowing the network to learn the optimal balance between the power spectrum path and the convolutional path. The addition of the power spectrum path introduces 256×256=65 53625625665536256$\times$ 256=65\,536256 × 256 = 65 536 extra parameters to the network, which is computational insignificant in comparison to both the size of the whole Psi-GAN network, as well as the alternative of adding a power spectrum term to the loss function and running a hyperparameter search in order to optimise its weighting.
@@
The addition of $L_{\mathrm{pixel}}$ to what is otherwise the standard WGAN-GP loss function was motivated by Piras et al. (2023), who found this term to aid in generating emulations with accurate statistics but ineffective in producing structure correlated with target $N$-body simulations.
@@ 3.3 Training
We train using the Adam optimiser (Kingma & Ba, 2014) with a base learning rate $\alpha=10^{-4}$, and decay parameters $(\beta_{1},\beta_{2})=(0.5,0.9)$. Following Heusel et al. (2017), we increase the learning rate for the critic by a factor of $f=3$ while using the base rate for the generator. We allow the model to train for an initial period of three epochs, after which we half the learning rate after every epoch where the validation loss increases. We also employ gradient clipping to clamp the magnitude of the gradients to a maximum value of 1 00010001\,0001 000. The gradient penalty term in the loss function should act to keep gradients close to unity, however there is a warm up period until it is able to have its intended effect. Clipping the gradients was found to be useful in avoiding overflow errors before the gradient penalty took effect.
@@ 3.4.5 Bispectrum
We measure the bispectra and reduced matter bispectra based on an estimator of the binned bispectrum (Coulton et al., 2019; Coulton & Spergel, 2019). Bispectra can be measured along different triangle configurations, and it is important to consider many configurations when using the bispectrum as a statistical tool in order to break degeneracies when inferring cosmological parameters (Berge et al., 2010). Therefore we measure the bispectra and reduced bispectra using multiple configurations of $(k_{1},k_{2})\in\{0.05,0.2,0.4,0.6\}~{}h~{}\mathrm{Mpc}^{-1}$, which span both regular configurations and squeezed bispectra configurations. In Section 4, we report Psi-GAN’s performance for only two of these configurations: $(k_{1},k_{2})=(0.4,0.4)~{}h~{}\mathrm{Mpc}^{-1}$ and $(k_{1},k_{2})=(0.4,0.6)~{}h~{}\mathrm{Mpc}^{-1}$, however our full set of tests show that Psi-GAN’s performance is similar over all configurations tested.
@@ 4 Results
Although Psi-GAN was trained with the goal of reproducing accurate statistics, we also see some correlations in structure between Psi-GAN emulations and $N$-body simulations. This emerges as the GAN framework aims to reproduce maps matching $N$-body simulations starting from correlated lognormal maps. However, for the applications we are interested in, we mainly care about summary statistics, and therefore choose to assess the model’s performance by how well it is able to reproduce those, as opposed to assessing any apparent structure correlation.
@@ 4.1 Randomised test set
We see agreement to within ${\sim}5$ per cent for all metrics, with the exception of the pixel counts at low values of $\mathrm{ln}(1+\delta)$. This is due to the baseline count for the $N$-body simulation being very low (${\sim}10^{2}$), and thus making the relative differences sensitive to small changes in pixel counts. We also believe that this issue is partially caused by the model’s architecture more easily modelling higher values of $\mathrm{ln}(1+\delta)$ due to the use of the GELU activation function (Hendrycks & Gimpel, 2016), which is more expressive at positive values.
@@ 5 Conclusions
The largest shortcoming of Psi-GAN is its slightly weaker performance in constraining the power spectrum when tasked with interpolating between cosmologies. In our tests, the power spectrum of samples generated by Psi-GAN showed less agreement with $N$-body simulations when interpolating between the cosmologies used in the Quijote simulations. An approximately ${\sim}50$ per cent greater coverage of cosmology space should be enough to reduce this shortcoming, however this would require significant resources to generate. Another potential method of improving this would be to pre-train the model architecture to be able to reconstruct lognormal random fields across an extensive dataset before fine-tuning the model to transform lognormal random field to more accurate emulations.
We used saliency mapping techniques to investigate further architectural improvements to Psi-GAN, which highlighted a slight weakness in capturing long-range correlations as well as a small issue with pixel-scale artefacting. Increasing the depth of the generator by another step (i.e. including an extra set of down-sample and up-sample blocks) should help Psi-GAN model long-range correlations better as the latent space will be more compressed and information will propagate more efficiently across the simulation box. Adding additional ConvNeXt blocks after the first convolution, and before the last convolution should aid in modelling small scales and reduce artefacting, as well as reduce the asymmetry between modelling negative and positive values of the matter density field as discussed in Section 4.1.